OpenStudy (anonymous):

What is the fourth term in the expansion of (x+2)6? 480x 168x4 160x3 120x3

OpenStudy (anonymous):

\[(x+2)^6=x^6+12 x^5+60 x^4+160 x^3+240 x^2+192 x+64 \]

OpenStudy (anonymous):

@robtobey can you explain?

OpenStudy (anonymous):

use the formula \[T _{r+1}=_{r}^{n}C(a)^{n-r}(b)^{r}\] for 4th term put r=3 in the formula where n=max power=6 a is first term (term independent of x ie 2) b is the second term ie x

OpenStudy (anonymous):

\[T _{3}+1=\frac{ 6 }{3 }C(a)^{3}(b)^{3}\]

OpenStudy (anonymous):

is that right so far?

OpenStudy (anonymous):

is the answer 160x3?

OpenStudy (anonymous):

no that's wrong it should be \[T _{4}=_{3}^{6}C (2)^{3}(x ^{3})\] where C is combination (6C3)

OpenStudy (anonymous):

\[T_{4}=2C(8)(x ^{3})\]

OpenStudy (anonymous):

is that closer to how you simplify it?

OpenStudy (anonymous):

no its not 6/3 C its 6C3 and is evaluated by typing 6 followed by pressing nCr key on your calculator to type C and then typing 3 it should give 20 as answer

OpenStudy (anonymous):

The rest is ok the answer should be 160x^3

OpenStudy (anonymous):

ohh okay, I get it. Thank you so much!

OpenStudy (anonymous):

np :)